# Chow functions for partially ordered sets

### v1

@@ 3.2. Definition of Chow functions
Let $\kappa$ be a $(P,\rho)$-kernel. We define the Chow function associated to $\kappa$, or $\kappa$-Chow function, as the element $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ defined by
@@
As a consequence of having defined $\overline{\kappa}_{ss}(x)$ as $-1$ for every $s\in P$, the minus sign appearing in the above definition guarantees that $\mathrm{H}_{ss}(x)=1$ for every $s\in P$. In the subsequent sections of this article we will focus our attention on a number of interesting examples of Chow functions. Notice that our definition of Chow functions as $-(\overline{\kappa})^{-1}$ is equivalent to either of the following properties:
@@
$\mathrm{H}_{st}(x)$ $=\sum_{s\leq w<t}\mathrm{H}_{sw}(x)\,\overline{\kappa}_{wt}(x),\qquad\text{for all $s<t$ in $P$}.$ (6)
@@
Let $\kappa$ be a $(P,\rho)$-kernel, and let $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ be the corresponding Chow function. Then, the following properties hold true:
@@
    In particular, if $\kappa$ is non-degenerate, we have that $\deg\mathrm{H}_{st}=\rho_{st}-1$ for every $s<t$.
@@
Hence, since $\mathrm{H}_{st}(x)=\overline{\kappa}_{st}(x)$, the last equation reduces to:
@@
Now, proceeding by induction of the size of the interval $[s,t]$, notice that every term in the sum on the right hand side of equation (7) has degree at most $\rho_{sw}-1+\rho_{wt}-1=\rho_{st}-2$. Therefore $[x^{\rho_{st}-1}]\mathrm{H}_{st}(x)=[x^{\rho_{st}-1}]\overline{\kappa}_{st}(x)=[x^{\rho_{st}}]\kappa_{st}(x)$. Now, to prove the second property, we multiply the formula of equation (7) by $x-1$ and subtract the polynomial $\kappa_{st}(x)$ from both sides, thus obtaining:
@@
where in (9) we used the induction hypothesis since $[w,t]$ is a strictly smaller interval, in (10) we used that $\kappa$ is a $(P,\rho)$-kernel, in (11) we used equation (8) again but changing the variable $x$ by $x^{-1}$ and $w$ by $u$, while in (13) we used again that $\kappa$ is a $(P,\rho)$-kernel. Note that it is possible to simplify the equality between the left-hand-side of (8) and the right-hand-side of (13):
@@
To conclude, observe that adding $\mathrm{H}_{st}(x)$ to both sides in equation (8), and combining the result with the last formula we obtained it is possible to see that $\mathrm{H}_{st}(x)=x^{\rho_{st}-1}\mathrm{H}_{st}(x^{-1})$, and the proof is complete. ∎
@@ 3.4. Non-negativity and unimodality of Chow functions
It is known that the right and left KLS functions arising from a $(P,\rho)$-kernel $\kappa$ can fail to be non-negative. Similarly, by our Example 3.5 a Chow function can fail to be non-negative too. As we will show now, there is a striking connection between the non-negativity of the KLS functions and the non-negativity and unimodality of the Chow function. We will prepare the proof with a few preliminary lemmas.
@@ 4.3. From characteristic Chow functions to Chow rings of matroids
The Chow ring of $\mathsf{M}$, denoted $\underline{\mathrm{CH}}(\mathsf{M})$ is defined as the quotient ring $S/(I+J)$. This is a graded ring, admitting a decomposition $\underline{\mathrm{CH}}(\mathsf{M})=\underline{\mathrm{CH}}^{0}(\mathsf{M})\oplus\cdots\oplus\underline{\mathrm{CH}}^{r-1}(\mathsf{M})$, where $r=\operatorname{rk}(\mathsf{M})$ is the rank of the matroid. The following was one of our main results in the prequel [FMSV24]. For the sake of completeness, we reformulate its proof adapted to the framework of the present paper.
@@
The preceding two formulas by Larson motivated us to search an analogous recursion for the right augmented Chow polynomial. The following result achieves so, but in a much less straightforward way. Furthermore, this is a non-trivial decomposition when viewed under the lens of the Hodge theory of matroids, because in that case it involves intersection cohomologies of truncations.
@@
$F_{P}(x)=Z_{P}(x)+x\sum_{t\neq\widehat{0}}F_{\operatorname{trunc}([\widehat{0},t])}(x)Z_{t\widehat{1}}(x).$
@@
$F_{P}(x)=Z_{P}(x)+x\sum_{\begin{subarray}{c}t\in P\\ t\neq\widehat{0}\end{subarray}}\mathrm{H}_{\operatorname{trunc}([\widehat{0},t])}(x)Z_{t\widehat{1}}(x),$
@@ 4.5. Unimodality, gamma-positivity, and flag $h$-vectors
In [FMSV24, Theorem 3.25] Ferroni, Matherne, Stevens, and Vecchi proved that when $P=\mathcal{L}(\mathsf{M})$ is the lattice of flats of the matroid $\mathsf{M}$, then the $\chi$-Chow polynomial and the left $\chi$-Chow polynomial are in fact $\gamma$-positive. The key ingredient in the proof is the semi-small decomposition for Chow rings and augmented Chow rings proved by Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22a].
Recently, based on a preliminary version of the theory developed in the present manuscript, Stump [Stu24] proved a more general result: if $P$ is a poset admitting an $R$-labelling, then the $\chi$-Chow polynomial and the left $\chi$-Chow polynomial are $\gamma$-positive. There is a strict chain of implications:
@@
Our goal now is to generalize Stump’s result even further, by proving that Cohen–Macaulayness is sufficient for the $\gamma$-positivity of $\chi$-Chow functions.
@@
The proof that we will provide is inherently technical, as one cannot a priori rely on any nice labelling for the poset. We will in fact prove a general relation between the flag $h$-vector and the $\chi$-Chow polynomial of $P$ (cf. Theorem 4.25), from where Theorem 4.20 will be an immediate corollary.
@@
As a consequence of the above statement and the non-negativity of the flag $h$-vector of a Cohen–Macaulay poset, we conclude the validity of Theorem 4.20. In an independent work, and simultaneously to the writing of the present paper, Liao [Lia24] proved a conjecture of Angarone, Nathanson, and Reiner [ANR23] on the equivariant $\gamma$-positivity of Chow rings of matroids, using techniques that are similar to our proof of Theorem 4.20.
@@
In [FMSV24, Conjecture 5.7] we conjectured that if $P$ is a geometric lattice, then $\mathrm{H}_{P}(x)$ interlaces $G_{P}(x)$. Experiments also suggest that this phenomenon may still be true for a Cohen–Macaulay poset. Under the assumption on $P$ being a geometric lattice, $F_{P}(x)$ appears to also be real-rooted and interlaced by $\mathrm{H}_{P}(x)$.
@@
    If $Z$ is $\gamma$-positive on all posets of $\mathcal{C}$, then so is $F$.
@@
$F_{st}(x)=Z_{st}(x)+x\sum_{w\neq s}F_{\operatorname{trunc}([s,w])}(x)\,Z_{wt}(x).$
Each summand $F_{\operatorname{trunc}([s,w])}(x)\,Z_{wt}(x)$ is a product of two symmetric polynomials. Assuming inductively that the first is unimodal (resp. $\gamma$-positive) and that the second is unimodal (resp. $\gamma$-positive), then so is their product. Furthermore, all of the summands are symmetric with center of symmetry $\frac{1}{2}(\rho(t)-\rho(s)-1)$. The factor $x$ preserves unimodality (resp. $\gamma$-positivity), and adding the unimodal (resp. $\gamma$-positive) term $Z_{st}(x)$, which has the correct center of symmetry, the proof follows. ∎
One of the main results of Ferroni, Matherne, Stevens, and Vecchi in [FMSV24, Theorem 4.7] establishes that when $P$ is a geometric lattice, the $Z$-polynomial arising from the characteristic function is $\gamma$-positive (unimodality was proved first via Hard Lefschetz by Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22b, Theorem 1.2(2)]). In particular, we obtain the following corollary to the above proposition.
@@
Note that the proof we indicated for geometric lattices relies on the heavy machinery of the singular Hodge theory of matroids, because the proof of the $\gamma$-positivity of the $Z$-polynomial uses the non-negativity of the Kazhdan–Lusztig polynomials of matroids [FMSV24, Remark 4.8]. In this case, the poset $P$ has right KLS polynomial equal to $f_{P}(x)=-3x+1$. This example shows that there is a huge leap between the behavior of $F$ and $Z$ for Cohen–Macaulay posets versus geometric lattices. For the sake of clarity, we have summarized these and further facts about characteristic Chow functions in Table 1.
@@ 5.1. Basics of Eulerian Chow polynomials
Let $n\in\mathbb{Z}_{\geq 0}$. From a $d$-dimensional convex polytope $\mathscr{P}\subseteq\mathbb{R}^{n}$ one can construct the poset $P$ of all of the faces of $\mathscr{P}$, where the order is given by the inclusion of sets. Note that $P$ is bounded and graded, and its rank equals $d$. The poset $P$ fulfills a number of important properties (see, e.g., [Zie95, Theorem 2.7]). One of the most important of such properties is that $P$ is Eulerian. This property, of fundamental relevance in algebraic combinatorics is defined as follows.
@@
An immediate conclusion from the last result is that the $\varepsilon$-Chow polynomial of an Eulerian poset $P$ is a non-negative combination of entries of the flag $h$-vector of $P$. In particular, if $P$ is Cohen–Macaulay then $\mathrm{H}_{P}(x)$ has non-negative coefficients. Even though there exist many Eulerian posets that are not Cohen–Macaulay (see, e.g., the poset on the right in Figure 1)., it is considerably hard to construct an Eulerian poset whose flag $h$-vector attains a negative entry. A subtle result by Bayer and Hetyei [BH01] shows that in fact all Eulerian posets of rank at most $6$ have a non-negative flag $h$-vector. However, they construct a very complicated example [BH01, Figure 2] in rank $7$ for which the flag $h$-vector attains a negative entry. Furthermore, modulo the Dehn–Sommerville relations, these flag $h$-vector attain a single negative entry, equal to $-1$. We refer to [Sta12, Solution to Exercise 193(b)] for a further discussion about that example.
@@ 5.2. Unimodality and $\gamma$-positivity
$\Psi_{P}(a,b)=\Phi_{P}(\mathbf{a}+\mathbf{b},\mathbf{ab}+\mathbf{ba}),$
@@ 6.1. A short recapitulation
Let $(W,S)$ be a Coxeter system and $T=\{wsw^{-1}\mid w\in W,s\in S\}$ the set of its reflections (in contrast with the set $S$ of simple reflections). The Bruhat order is the poset defined on $W$ with relations $u\leq v$ whenever there exist $w_{0},\ldots,w_{n}$ such that $w_{0}=u$, $w_{n}=v$ and $w_{i}^{-1}w_{i+1}=t\in T$ for all $i=1,\ldots,n-1$. This poset is graded, where the rank function is given by $\rho(w)=\ell(w)$, where $\ell(w)$ is the length of any reduced word equal to $w$. Furthermore, Bruhat intervals are Eulerian and shellable, and hence Gorenstein* (see [BB05, Proposition 2.7.5]).
@@ Acknowledgements
The authors wish to thank Louis Billera, Tom Braden, Petter Brändén, Chris Eur, Eva Feichtner, Matt Larson, Mateusz Michalek, Nicholas Proudfoot, Matthew Stevens, and Botong Wang for insightful conversations. Part of these results were obtained while the authors were visiting the Oberwolfach Research Institute of Mathematics MFO, as part of the workshop “Arrangements, matroids, and logarithmic vector fields”. We are grateful to MFO and the organizers of the workshop.

### v2

@@ 3.2. Definition of Chow functions
Let $\kappa$ be a $(P,\rho)$-kernel. We define the Chow function associated to $\kappa$, or the $\kappa$-Chow function, as the element $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ defined by
@@
As a consequence of having defined $\overline{\kappa}_{ss}(x)$ as $-1$ for every $s\in P$, the minus sign appearing in the above definition guarantees that $\mathrm{H}_{ss}(x)=1$ for every $s\in P$. Furthermore, we note that Definition 3.3 is equivalent to either of the following properties. For all $s<t$ in $P$,
@@
$\mathrm{H}_{st}(x)$ $=\sum_{s\leq w<t}\mathrm{H}_{sw}(x)\,\overline{\kappa}_{wt}(x).$ (6)
In the subsequent sections of this article we will focus our attention on a number of interesting examples of Chow functions.
@@
Let $\kappa$ be a $(P,\rho)$-kernel, and let $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ be the corresponding Chow function. The following properties hold.
@@
    In particular, if $\kappa$ is non-degenerate, then $\deg\mathrm{H}_{st}(x)=\rho_{st}-1$ for every $s<t$.
@@
Hence, since $\mathrm{H}_{st}(x)=\overline{\kappa}_{st}(x)$, the last equation reduces to
@@
Now, proceeding by induction on the size of the interval $[s,t]$, every term in the sum on the right hand side of equation (7) has degree at most $\rho_{sw}-1+\rho_{wt}-1=\rho_{st}-2$. Therefore $[x^{\rho_{st}-1}]\mathrm{H}_{st}(x)=[x^{\rho_{st}-1}]\overline{\kappa}_{st}(x)=[x^{\rho_{st}}]\kappa_{st}(x)$. Now, to prove the second property, we multiply the formula of equation (7) by $x-1$ and subtract the polynomial $\kappa_{st}(x)$ from both sides:
@@
where in (9) we used the induction hypothesis since $[w,t]$ is a strictly smaller interval, in (10) we used that $\kappa$ is a $(P,\rho)$-kernel, in (11) we used equation (8) again but changing the variable $x$ by $x^{-1}$ and $w$ by $u$, while in (13) we used again that $\kappa$ is a $(P,\rho)$-kernel. Note that it is possible to simplify the equality between the left hand side of (8) and the right hand side of (13):
@@
To conclude, observe that adding $\mathrm{H}_{st}(x)$ to both sides in equation (8) and combining the result with the last formula we obtained, it is possible to see that $\mathrm{H}_{st}(x)=x^{\rho_{st}-1}\mathrm{H}_{st}(x^{-1})$, and the proof is complete. ∎
@@ 3.4. Non-negativity and unimodality of Chow functions
It is known that the right and left KLS functions arising from a $(P,\rho)$-kernel $\kappa$ can fail to be non-negative. Similarly, by our Example 3.5 a Chow function can fail to be non-negative too. As we will show now, there is a striking connection between the non-negativity of the KLS functions and the non-negativity and unimodality of the Chow function. The proof will need the following preparatory lemma.
@@ 4.3. From characteristic Chow functions to Chow rings of matroids
The Chow ring of $\mathsf{M}$, denoted $\underline{\mathrm{CH}}(\mathsf{M})$ is defined as the quotient ring $S/(I+J)$. This is a graded ring, admitting a decomposition $\underline{\mathrm{CH}}(\mathsf{M})=\underline{\mathrm{CH}}^{0}(\mathsf{M})\oplus\cdots\oplus\underline{\mathrm{CH}}^{r-1}(\mathsf{M})$, where $r=\operatorname{rk}(\mathsf{M})$ is the rank of the matroid. The following was one of the main results of [FMSV24]. For the sake of completeness, we reformulate its proof adapted to the framework of the present paper.
@@
The preceding two formulas by Larson motivated us to search an analogous recursion for the right augmented Chow polynomial. The following result achieves so, but in a much less straightforward way. Furthermore, this is a non-trivial decomposition when viewed under the lens of the Hodge theory of matroids (see the discussion at the end of Section 4.5).
@@
$F_{P}(x)=Z_{P}(x)+x\sum_{\begin{subarray}{c}t\in P\\ \rho(t)>1\end{subarray}}\mathrm{H}_{\operatorname{trunc}([\widehat{0},t])}(x)Z_{t\widehat{1}}(x).$
@@
$F_{P}(x)=Z_{P}(x)+x\sum_{\begin{subarray}{c}t\in P\\ \rho(t)>1\end{subarray}}\mathrm{H}_{\operatorname{trunc}([\widehat{0},t])}(x)Z_{t\widehat{1}}(x),$
@@ 4.5. Unimodality, gamma-positivity, and flag $h$-vectors
When $P=\mathcal{L}(\mathsf{M})$ is the lattice of flats of the matroid $\mathsf{M}$, then the $\chi$-Chow polynomial and the left $\chi$-Chow polynomial are in fact $\gamma$-positive (see [FMSV24, Theorem 3.25]). The key ingredient in the proof is the semi-small decomposition for Chow rings and augmented Chow rings proved by Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22a].
Recently, using notions in a preliminary version of the theory developed in the present manuscript, Stump [Stu24] proved a more general result: if $P$ is a poset admitting an $R$-labelling, then the $\chi$-Chow polynomial and the left $\chi$-Chow polynomial are $\gamma$-positive. There is a strict chain of implications:
@@
The proof that we will provide is inherently technical, as one cannot a priori rely on any nice labelling for the poset. We will in fact prove a general relation between the flag $h$-vector and the $\chi$-Chow polynomial of $P$ (cf. Theorem 4.25), from where Theorem 4.20 and Stump’s result will be immediate corollaries.
@@
As a consequence of the above statement and the non-negativity of the flag $h$-vector of a Cohen–Macaulay poset, we conclude the validity of Theorem 4.20. Observe that this also proves Stump’s result, because posets with an $R$-labelling have non-negative flag $h$-vectors. However, the classes of Cohen–Macaulay posets and posets with $R$-labellings do not compare to each other.
In an independent work, and simultaneously to the writing of the present paper, Liao [Lia24] proved a conjecture of Angarone, Nathanson, and Reiner [ANR23] on the equivariant $\gamma$-positivity of Chow rings of matroids, using techniques that are similar to our proof of Theorem 4.20.
@@
We also point out that the above conjecture bears a resemblance to the Neggers–Stanley conjecture for which there are known counterexamples by Stembridge [Ste07] (see also the previous work by Brändén [Brä04]). However, we have verified with a computer that none of the lattices of ideals of Stembridge’s counterexamples to the Neggers–Stanley conjecture disproves Conjecture 4.26.
@@
In [FMSV24, Conjecture 5.7], $\mathrm{H}_{P}(x)$ is conjectured to interlace $G_{P}(x)$, whenever $P$ is a geometric lattice. Experiments also suggest that this phenomenon may still be true for a Cohen–Macaulay poset. Under the assumption on $P$ being a geometric lattice, $F_{P}(x)$ appears to also be real-rooted and interlaced by $\mathrm{H}_{P}(x)$.
@@
    If $\mathrm{H}$ and $Z$ are $\gamma$-positive on all posets of $\mathcal{C}$, then so is $F$.
@@
$F_{st}(x)=Z_{st}(x)+x\sum_{w\neq s}\mathrm{H}_{\operatorname{trunc}([s,w])}(x)\,Z_{wt}(x).$
Each summand $\mathrm{H}_{\operatorname{trunc}([s,w])}(x)\,Z_{wt}(x)$ is a product of two symmetric polynomials. Since both polynomials are always unimodal (resp. $\gamma$-positive), then so is their product. Furthermore, all of the summands are symmetric with center of symmetry $\frac{1}{2}(\rho(t)-\rho(s)-2)$. The factor $x$ preserves unimodality (resp. $\gamma$-positivity), and shifts the center of symmetry to $\frac{1}{2}(\rho(t)-\rho(s))$. By adding the unimodal (resp. $\gamma$-positive) term $Z_{st}(x)$, which has the correct center of symmetry, the proof follows. ∎
One of the main results of [FMSV24, Theorem 4.7] establishes that when $P$ is a geometric lattice, the $Z$-polynomial arising from the characteristic function is $\gamma$-positive (unimodality was proved first via Hard Lefschetz by Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22b, Theorem 1.2(2)]), proving a conjecture in [FNV23]. In particular, we obtain the following corollary to the above proposition.
@@
Note that the proof of $\gamma$-positivity we have for geometric lattices relies on the heavy machinery of the singular Hodge theory of matroids, because the proof of the $\gamma$-positivity of the $Z$-polynomial uses the non-negativity of the Kazhdan–Lusztig polynomials of matroids [FMSV24, Remark 4.8]. In this case, the poset $P$ has right KLS polynomial equal to $f_{P}(x)=-3x+1$. This example shows that there is a huge leap between the behavior of $F$ and $Z$ for Cohen–Macaulay posets versus geometric lattices. For the sake of clarity, we have summarized these and further facts about characteristic Chow functions in Table 1.
@@ 5.1. Basics of Eulerian Chow polynomials
Let $n\in\mathbb{Z}_{\geq 0}$. From a $d$-dimensional convex polytope $\mathscr{P}\subseteq\mathbb{R}^{n}$ one can construct the poset $P$ of all of the faces of $\mathscr{P}$, where the order is given by inclusion. Note that $P$ is bounded and graded, and its rank equals $d$. The poset $P$ fulfills a number of important properties (see, e.g., [Zie95, Theorem 2.7]). One of the most important of such properties is that $P$ is Eulerian. This property, of fundamental relevance in algebraic combinatorics, is defined as follows.
@@
An immediate conclusion from the last result is that the $\varepsilon$-Chow polynomial of an Eulerian poset $P$ is a non-negative combination of entries of the flag $h$-vector of $P$. In particular, if $P$ is Cohen–Macaulay then $\mathrm{H}_{P}(x)$ has non-negative coefficients. Even though there exist many Eulerian posets that are not Cohen–Macaulay (see, e.g., the poset on the right in Figure 1), it is considerably hard to construct an Eulerian poset whose flag $h$-vector attains a negative entry. A subtle result by Bayer and Hetyei [BH01] shows that in fact all Eulerian posets of rank at most $6$ have a non-negative flag $h$-vector. However, they construct a very complicated example [BH01, Figure 2] in rank $7$ for which the flag $h$-vector attains a negative entry. Furthermore, modulo the Dehn–Sommerville relations, these flag $h$-vector attain a single negative entry, equal to $-1$. We refer to [Sta12, Solution to Exercise 193(b)] for further discussion about that example.
@@ 5.2. Unimodality and $\gamma$-positivity
$\Psi_{P}(\mathbf{a},\mathbf{b})=\Phi_{P}(\mathbf{a}+\mathbf{b},\mathbf{ab}+\mathbf{ba}),$
@@ 6.1. A short recapitulation
Let $(W,S)$ be a Coxeter system and $T=\{wsw^{-1}\mid w\in W,s\in S\}$ the set of its reflections (in contrast with the set $S$ of simple reflections). The Bruhat order is the poset defined on $W$ with relations $u\leq v$ whenever there exist $w_{0},\ldots,w_{n}$ such that $w_{0}=u$, $w_{n}=v$ and $w_{i}^{-1}w_{i+1}=t\in T$ for all $i=1,\ldots,n-1$. This poset is graded, with rank function given by $\rho(w)=\ell(w)$, where $\ell(w)$ is the length of any reduced word equal to $w$. Furthermore, Bruhat intervals are Eulerian and shellable, and hence Gorenstein* (see [BB05, Proposition 2.7.5]).
@@ Acknowledgements
The authors wish to thank Christos Athanasiadis, Louis Billera, Tom Braden, Petter Brändén, Chris Eur, Eva Feichtner, June Huh, Matt Larson, Mateusz Michalek, Nicholas Proudfoot, Matthew Stevens, and Botong Wang for insightful conversations. Part of these results were obtained while the authors were visiting the Oberwolfach Research Institute of Mathematics MFO, as part of the workshop “Arrangements, matroids, and logarithmic vector fields”. We are grateful to MFO and the organizers of the workshop.